# On a generalisation of Cameron’s base size conjecture

### v2

@@ 9.4. Graph automorphisms
Now assume that $q=2$. Here, the value $g(x,q)$ can be obtained by Proposition 8.2. For parabolic subgroups, the unipotent $\bar{G}_{\sigma}$-classes and the tables of Foulkes functions are given to us again by Lübeck in [64], so we can compute precise fixed point ratios for involutions in parabolic actions using GAP.
The odd order semisimple classes are listed in [25, Table 9]. From this table we can recover the pair $(J,[w])$ corresponding to each semisimple $\bar{G}_{\sigma}$-class via [39]. More specifically, the centraliser type for each class is listed in [25, Table 9], so we can immediately recover $J$ by consulting [39], and by inspecting the possible centraliser orders for this particular choice of $J$ we can then recover $w$. For example, assume that $\epsilon=+$ and let us look at the fourth column in [25, Table 9]. In this case our element $y$ has centraliser type $A_{3}T_{3}$. We then turn to [39] and we find that $J=\{\alpha_{0},\alpha_{2},\alpha_{4}\}$ and by computing all possible centraliser orders we deduce that $w=a_{1}a_{17}a_{18}$. In this way we can compute precise fixed point ratios for semisimple elements, as in the case $q\geqslant 3$.
@@
If $\epsilon=$-, then we can show that $\widehat{Q}(G,\tau)<1$ for any $6$-tuple $\tau$ of $G$ using almost the same strategy as in the case $q\geqslant 3$. The only difference is that we only considering prime order elements when finding upper bounds for $\alpha$ and $\beta$. In particular, (13) still holds, where, as mentioned above, $g(x,q)$ is obtained via Proposition 8.2. Then using (15), we find that $\alpha\leqslant 0.1$ and by using Lübeck’s data on unipotent elements we find that $\beta\leqslant 0.055$. Finally, the contribution from involutory graph automorphisms is no more than $0.125$, and hence $\widehat{Q}(G,\tau)<1$.
The case $\epsilon=+$ requires special attention. As pointed out in [25], if $G_{0}=E_{6}(2)$ and $\tau$ is a conjugate $6$-tuple containing copies of $P_{1}$ or $P_{6}$, then $\widehat{Q}(G,\tau)>1$. Therefore, it is not possible to prove the claim using the same methods as in the case $q\geqslant 3$, so we adopt a slightly different approach. We find that it is possible to construct $P_{1}$ and $P_{6}$ in Magma, and using random search, one can show that every $6$-tuple of $G$ only containing copies of $P_{1}$ and $P_{6}$ is regular. We may therefore assume that $\tau$ contains at least one entry not isomorphic to $P_{1}$ or $P_{6}$.
@@
We can now estimate $\alpha^{\prime},\beta^{\prime},\gamma^{\prime}$ from above using the methods described above, and we find $\alpha^{\prime}\leqslant 0.053$, $\beta^{\prime}\leqslant 0.6$, and $\gamma^{\prime}\leqslant 0.125$, so $\widehat{Q}(G,\tau)<1$, and the proof is complete. ∎
@@
We find in [65] that the total number of semisimple classes in $G_{\sigma}$ is no more than $q^{7}+q^{4}$ and unless $C_{G}(x)$ has an $E_{6},D_{6}$, or $E_{7}$ factor, it suffices to take $C_{i}=q^{7}+q^{4}$. In the final few cases we can estimate $c_{i}$ by inspecting [65] and we give a bound $C_{i}$ in Table 19. In particular, we record the type of $C_{\bar{G}}(x_{i})$, the set $J$ and $w$ and an upper bound on the number of classes of such type. Recall that we use the standard Bourbaki labelling [11] when recording the set $J$ in this paper, instead of the notation used in [39]. Now using Magma, we can compute the expression
@@
We use the same approach we used in the proof of Proposition 10.7 when $q\geqslant 3$. In particular, (13) holds in this case as well, where $g(x,q)$ is obtained this time via Proposition 8.4 for each choice of prime order element $x\in G$, so we can estimate $\alpha,\beta$, and $\gamma$, as defined in (12) to show that $\widehat{Q}(G,\tau)<1$.
First let us consider semisimple elements. In this case the pairs $(J,[w])$ encoding the possible centraliser types of semisimple elements can be found in [3, Table 23] and the number of classes of each centraliser type can be found in [65] as usual, so we can compute a bound on $\alpha$ via (15) and we find that $\alpha<q^{-1}$.
@@
${\rm fpr}(x,G/H)\leqslant\frac{|F_{4}^{\epsilon}(q^{1/r})\,:\,\bar{H}^{\epsilon}(q^{1/r})|}{|F_{4}(q)\,:\,\bar{H}(q)|}.$
In this way, we find that $f(x,q)\leqslant q^{-6}$ if $r=2$, whilst $f(x,q)$ is achieved when $\bar{H}=P_{1}$ if $r\geqslant 3$. Noting that $|x^{G}|<2q^{26}$ if $x$ is an involutory field or graph-field automorphism, we find
@@
Now suppose that $x$ is unipotent. In this case inspecting [59, Table 22.2.1] tells us that the smallest class of elements in $\mathcal{L}_{2}$ is $A_{3}A_{2}A_{1}$ and the largest is $D_{5}$. We therefore find
@@
We now show that $f(x,q)\leqslant c_{i}$ for all $i\in\{1,\ldots,5\}$. Fix an $i\in\{1,\ldots,5\}$, a prime order element $x\in G$ such that $a_{i}\leqslant{\rm dim}\,x^{\bar{G}}<b_{i}$ and $H\in\mathcal{M}_{G}$. If $i=5$, then [51, Theorem 1] implies that ${\rm fpr}(x,G/H)\leqslant 2q^{-24}=c_{5}$, so we may assume that $i\leqslant 4$. First assume that $H$ is non-parabolic. If $|H|\leqslant q^{88}$, then the claim holds from the claim above as ${\rm fpr}(x,G/H)\leqslant|H|/|x^{G}|\leqslant c_{i}$.
Now assume that $|H|>q^{88}$. The possibilities for $H$ are described in [25, 4.1] and one can check that ${\rm fpr}(x,G/H)\leqslant c_{i}$ by inspecting the proofs of [25, 4.3–4.5].
@@
It now remains to show that the number of prime order elements $x\in G$ with $a_{i}\leqslant{\rm dim}\,x^{\bar{G}}<b_{i}$ is bounded above by $d_{i}$ for all $i\in\{1,\ldots,5\}$. We treat each possibility of $i$ in turn.
@@
Note that if $182\leqslant{\rm dim}\,x^{\bar{G}}<202$, then $x$ is either semisimple or unipotent, or a field automorphism of order $5$. It follows from [65] that $G$ has fewer than $2q^{8}$ semisimple classes and we verify using [64] that there are exactly 24 unipotent classes with $182\leqslant{\rm dim}\,x^{\bar{G}}<202$. Finally, there are no more than 4 classes of field automorphisms of order $5$, and $|x^{G_{0}}|<q^{202}$, as shown above, so we get
@@
Here $154\leqslant{\rm dim}\,x^{\bar{G}}<182$, and so $x$ is either semisimple or unipotent, or a field automorphism of order $3$. By inspecting [65] we deduce that there are at most $2q^{2}+3q+6$ semisimple classes with $154\leqslant{\rm dim}\,x^{\bar{G}}<182$ and it follows from [64] that there are exactly 11 unipotent classes whose dimension is in this range. Finally there are no more than $2$ classes of involutory field automorphisms. It is also clear that $|x^{G}|<q^{182}$, and so we get

### v3

@@ 9.4. Graph automorphisms
Now assume that $q=2$. Here, the value $g(x,q)$ can be obtained by Proposition 8.2. For parabolic subgroups, the unipotent $\bar{G}_{\sigma}$-involutions and the tables of Foulkes functions are are determined by Lübeck in [65], so we can compute precise fixed point ratios for involutions in parabolic actions using GAP.
The odd order semisimple classes are listed in [26, Table 9]. From this table we can recover the pair $(J,[w])$ corresponding to each semisimple $\bar{G}_{\sigma}$-class via [40]. More specifically, the centraliser type for each class is listed in [26, Table 9], so we can immediately recover $J$ by consulting [40], and by inspecting the possible centraliser orders for this particular choice of $J$, we can then recover $w$.
For example, assume that $\epsilon=+$ and let us look at the fourth row in [26, Table 9]. In this case our element $y$ has centraliser type $A_{3}T_{3}$. We then turn to [40] and we find that $J=\{\alpha_{0},\alpha_{2},\alpha_{4}\}$ and by computing all possible centraliser orders we deduce that $w=a_{1}a_{17}a_{18}$. In this way we can compute precise fixed point ratios for all prime order semisimple elements, as in the case $q\geqslant 3$.
@@
If $\epsilon=$-, then we can show that $\widehat{Q}(G,\tau)<1$ for any $6$-tuple $\tau$ of $G$ using almost the same strategy as in the case $q\geqslant 3$. The only difference is that we are only considering prime order elements when finding upper bounds for $\alpha$ and $\beta$. In particular, (15) still holds, where, as mentioned above, $g(x,q)$ is obtained via Proposition 8.2. Then using (17), we find that $\alpha\leqslant 0.1$ and by using Lübeck’s data on unipotent elements we find that $\beta\leqslant 0.055$. Finally, the contribution from involutory graph automorphisms is no more than $0.125$, and hence $\widehat{Q}(G,\tau)<1$.
The case $\epsilon=+$ requires special attention. As pointed out in [26], if $G_{0}=E_{6}(2)$ and $\tau$ is a conjugate $6$-tuple containing copies of $P_{1}$ or $P_{6}$, then $\widehat{Q}(G,\tau)>1$. Therefore, it is not possible to prove the claim using the same methods as in the case $q\geqslant 3$, so we adopt a slightly different approach. We find that it is possible to construct $P_{1}$ and $P_{6}$ in Magma. More specifically, we can construct a $P_{1}$-parabolic using the following code:
G:=AutomorphismGroupSimpleGroup("E6",2);
T:=Socle(G);
H:=Stabilizer(T,1);
and then noting that $P_{1}$ and $P_{6}$ are $G_{0}.2$-conjugate but not $G_{0}$-conjugate, we can construct $P_{6}$ using random search. Then again using random search, one can show that every $6$-tuple of $G$ only containing copies of $P_{1}$ and $P_{6}$ is regular. We may therefore assume that $\tau$ contains at least one component that is not isomorphic to $P_{1}$ or $P_{6}$.
@@
We can now estimate $\alpha^{\prime},\beta^{\prime},\gamma^{\prime}$ from above using the methods described above, and we find that $\alpha^{\prime}\leqslant 0.053$, $\beta^{\prime}\leqslant 0.6$, and $\gamma^{\prime}\leqslant 0.125$, so $\widehat{Q}(G,\tau)<1$ and the proof is complete. ∎
@@
We find in [66] that the total number of semisimple classes in $\bar{G}_{\sigma}$ is no more than $q^{7}+q^{4}$ and unless $C_{G}(x)$ has an $E_{6},D_{6}$, or $E_{7}$ factor, it suffices to take $C_{i}=q^{7}+q^{4}$. In the final few cases we can estimate $c_{i}$ by inspecting [66] and we give a bound $C_{i}$ in Table 21. In particular, we record the type of $C_{\bar{G}}(x_{i})$, the set $J$ and $w$ and an upper bound on the number of classes of such type. Recall that we use the standard Bourbaki labelling [11] when recording the set $J$ in this paper, instead of the notation used in [40]. Now using Magma, we can compute the expression
@@
We use the same approach we did in the proof of Proposition 10.7 when $q\geqslant 3$. In particular, (15) holds in this case as well, where $g(x,q)$ is obtained this time via Proposition 8.4 for each choice of prime order element $x\in G$, so we can estimate $\alpha,\beta$, and $\gamma$, as defined in (14) to show that $\widehat{Q}(G,\tau)<1$.
First let us consider semisimple elements. In this case, the pairs $(J,[w])$ encoding the possible centraliser types of semisimple elements can be found in [3, Table 23] and the number of classes of each centraliser type can be found in [66] as usual, so we can compute a bound on $\alpha$ via (17), and we find that $\alpha<q^{-1}$.
@@
${\rm fpr}(x,G/H)\leqslant\frac{|F_{4}^{\epsilon}(q^{1/r})\,:\,\bar{H}^{\epsilon}(q^{1/r})|}{|F_{4}(q)\,:\,\bar{H}(q)|},$
where $\bar{H}(q)$ and $\bar{H}(q^{1/r})$ are defined as in Section 9.3. In this way, we find that $f(x,q)\leqslant q^{-6}$ if $r=2$, whilst $f(x,q)$ is achieved when $\bar{H}=P_{1}$ if $r\geqslant 3$. Noting that $|x^{G}|<2q^{26}$ if $x$ is an involutory field or graph-field automorphism, we find
@@
Now suppose that $x$ is unipotent. In this case inspecting [60, Table 22.2.1] tells us that the smallest class of elements in $\mathcal{L}_{2}$ is $A_{3}A_{2}A_{1}$ and the largest is $D_{5}$. We therefore find
@@
We now show that $f(x,q)\leqslant c_{i}$ for all $i\in\{1,\ldots,5\}$. Fix an $i\in\{1,\ldots,5\}$, a prime order element $x\in G$ such that $a_{i}\leqslant{\rm dim}\,x^{\bar{G}}<b_{i}$ and $H\in\mathcal{M}_{G}$. If $i=5$, then [52, Theorem 1] implies that ${\rm fpr}(x,G/H)\leqslant 2q^{-24}=c_{5}$, so we may assume that $i\leqslant 4$. First assume that $H$ is non-parabolic. If $|H|\leqslant q^{88}$, then the claim follows from the bounds on $|x^{G}|$ above, as ${\rm fpr}(x,G/H)\leqslant|H|/|x^{G}|\leqslant c_{i}$.
Now assume that $|H|>q^{88}$. The possibilities for $H$ are described in [26, 4.2] and one can check that ${\rm fpr}(x,G/H)\leqslant c_{i}$ by inspecting the proofs of [26, 4.4–4.6].
@@
where $\bar{H}(q)$ and $\bar{H}(q^{1/r})$ are defined as in Section 9.3. It now remains to show that the number of prime order elements $x\in G$ with $a_{i}\leqslant{\rm dim}\,x^{\bar{G}}<b_{i}$ is bounded above by $d_{i}$ for all $i\in\{1,\ldots,5\}$. We treat each possibility of $i$ in turn.
@@
Here, we either have $x\in\bar{G}_{\sigma}$ or $x$ is a field automorphism of order $5$. It follows from [66] that $G$ has fewer than $2q^{8}$ semisimple classes and we verify using [65] that there are exactly 24 unipotent classes with $182\leqslant{\rm dim}\,x^{\bar{G}}<202$. Finally, there are no more than 4 classes of field automorphisms of order $5$, and $|x^{G_{0}}|<q^{202}$, as shown above, so we get
@@
Here, we either have $x\in\bar{G}_{\sigma}$ or $x$ is a field automorphism of order $3$. By inspecting [66], we deduce that there are at most $2q^{2}+3q+6$ semisimple classes with $154\leqslant{\rm dim}\,x^{\bar{G}}<182$ and it follows from [65] that there are exactly 11 unipotent classes whose dimension is in this range. Finally, there are no more than $2$ classes of involutory field automorphisms. We’ve also shown above that $|x^{G}|<q^{182}$, and so we get